OpenStudy (anonymous):

HELP! You are given two points (1,3) and (2,4) which equation best fits these points? Y=x-2 Y=2x+1 Y=12x+4 Y=x+2

OpenStudy (anonymous):

It is the first one, Y=x-2 because if you put the points into the equation you would have 1=3-2 and 2=4-2

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

Yes! I understand exactly now. Thank you for the help!

OpenStudy (anonymous):

No problem